OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form. (Points : 1)

OpenStudy (anonymous):

This is on my test that I am having a hard time on with

OpenStudy (anonymous):

slope intercept form. y = mx+b

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

can you help me with this one too, I don't like math at all. If you are finding the line perpendicular to x + 3y = 5, what is the slope of the line? (Points : 1) -3/5 1/3 3 -3

OpenStudy (anonymous):

put it in to y = mx+b form first.. m is the slope.. so to find perpendicular line, find the slope's reciprocal with the opposite sign

OpenStudy (anonymous):

okay thanks!!!

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I have to go I have test I have to do :(

OpenStudy (anonymous):

did you get the answers though?